a charge q is projected from a large distance with speed v towards a fixed charged particle Q. It approaches Q upto a distance r and then returns. If closest is r/2, then projected velocity should be

Grade:12

1 Answers

Vikas TU
7 years ago
At large distance that is the distance approaching to infinity the potential energy assumes to be zero.
Therefore, the energy theorem proves here to be:
0.5mv^2 + 0 = KQq/r.................(1)
for closest distance r/2,
0.5mv0^2 = KQq/(r/2) => 2KQq/r...................(2)
Dividing bot the eqns. we get,
(v/v0)^2 = ½
v/v0 = 1/root(2)
v0 = v*root(2) m/s